OpenStudy (anonymous):

Mrs. Computer is 3 times older than her daughter, Mousy. The sum of their ages is 52. How old is Mousy? Mousy is ____ years old. (Be sure to only type Mousy’s age in the answer blank.)

OpenStudy (shane_b):

Let x = Mrs. Computer and y = Mousy: x = 3y x+y=52 Solve it the same way as before...

OpenStudy (anonymous):

3x=y and x+y =52

OpenStudy (anonymous):

x=52/4 x = 13

OpenStudy (anonymous):

3x , 3(13)=39

OpenStudy (anonymous):

y=52-39, y=13

OpenStudy (anonymous):

so answer is 13

OpenStudy (anonymous):

Mrs. computter has 39 years old and her daughter is 13 years old.

OpenStudy (anonymous):

Thanks!